# Importance sampling for Sobol’ indices estimation

### v1

@@ 4 Numerical Experiments
4.1 Analytical Illustration of Variance Reduction via IS
In this section, we illustrate the effectiveness of an IS transformation in reducing the variance $\eta_{u}$ estimators. As a benchmark example, we consider the classical Sobol’ g-function [8]:
@@ 4.1 Analytical Illustration of Variance Reduction via IS
where $X_{i}\sim\mathcal{U}[0,1]$ are assumed to be independent inputs. This function is widely used in sensitivity analysis due to its non-linear, non-monotonic nature and the availability of analytical Sobol’ indices. We consider a 3-dimensional case $(k=3)$ with $a_{i}=i$, and focus on estimating the second-order Sobol’ index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$. The third input $X_{3}$ is treated as an uncontrolled variable $W$. We consider the case A in (3.3), where $\bar{u}=\emptyset$. In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically, including the optimal sampling density. Specifically, the optimal density $q^{\star}$ to estimate $\eta_{12}$ is given by
@@
We first plot the shape of $q^{*}$ over $[0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
@@
Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte-Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
(a) Optimal density $q^{\star}(x_{1},x_{2})$
@@
Figure 1: Left: optimal sampling density $q^{\star}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{\star}$.
Parametric optimization
@@
where the shape parameters $\alpha_{i},\beta_{i}$ are positive real numbers, i.e., $\alpha_{i},\beta_{i}\in\mathbb{R}_{+}^{*}$ for $i=1,2$. We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform distribution at $(\alpha,\beta)=(1,1)$. The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimizes the asymptotic variance, namely
@@ 4.2 Distributional Sensitivity via Reverse Importance Sampling
We study the sensitivity of the quantity
$\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m_{u}^{2}(\mathbf{X}_{u})\right],$
with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$. This falls into the framework of reverse importance sampling; this method, originally referred to as importance sampling sensitivity analysis [6], is now often described as a reverse importance sampling mechanism [27] and exploits the fact that a single set of observations can be reused to estimate responses under multiple reference distributions. Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [5].
The rank estimator [41], based on the sequential correlation of outputs, takes as input a sequence $Y^{(i)}$ sorted with respect to the associated inputs $\mathbf{X}_{u}^{(i)}$, and converges to $\eta_{u}$ via the average of consecutive products. While initially defined under the distribution $p$, this estimator can be naturally extended to estimate $\eta_{u}(\theta)$ under any distribution $p_{\theta}$, using the results from Section 3.1, the convergence to $\eta_{u}(\theta)$ is guaranteed.
For each $\theta$, we define the reweighted output
$Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\cdot\frac{p_{\theta,\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}\cdot Y,$ (19)
so that the rank-based estimator becomes
$\hat{\eta}^{\text{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)}.$ (20)
@@
[…]
@@
With the rank estimator in eq(4.3), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$. Two perturbation strategies are considered: (i) marginal perturbation, where only the pair $\theta_{j}=(\alpha_{j},\beta_{j})$ associated with a single input $X_{j}$ is varied while the remaining marginals remain uniform; and (ii) global perturbation, where several or all input marginals are varied simultaneously.
To keep the presentation simple, we focus on marginal perturbations. Specifically, we assess the distributional sensitivity of the contrast $\eta_{u}$ with $u=\{4\}$ by varying $\theta_{j}$ for one input variable at a time, keeping all others unchanged. This yields a family of perturbed estimates $\eta_{4}(\theta_{j})$, each revealing how the index $\eta_{4}$ responds to changes in the distribution of a specific input $X_{j}$, for $j=1,\dots,4$. All computations are based on the same reference dataset.
@@
These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, with markedly different behaviors across the four inputs. Overall, these confirm that the proposed given-data IS framework offers a powerful and cost-effective mean to perform detailed distributional local sensitivity analyses of Sobol’ indices.
@@ 5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution different from the reference one. In particular, we show that by choosing optimal densities, it is possible to achieve a lower bound on the asymptotic variance than the efficiency bound established in [19]. In some special cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. We also show that combining our method with the rank-based estimator [41] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [27]. Our approach can be naturally extended to other contexts, such as the given-model setting where a computational budget is available to evaluate the black-box model. In such cases, we expect to approximate the optimal densities presented in our framework and achieve significant variance reduction. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
Appendix A Lemma 1 and other derivations
Proof of Lemma 1
We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
@@ Appendix A Lemma 1 and other derivations
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}),$
@@
$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m_{u}^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m_{u}^{2}(\mathbf{X}_{u})]$
Derivation of the optimal variance under $p_{u}$ in (3.7)
We start from the expression
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$
Using the last identity we deduce
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}=w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u}),\quad\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}=w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u}).$
For the second term $\mathbb{E}_{q_{\bar{u}}}[Z_{u}^{2}\mid\mathbf{X}_{u}]$, we write
$\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]$ $=\mathbb{E}_{q}\left[\left(\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y\right)^{2}\middle|\mathbf{X}_{u}\right]$
$=w_{u}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
Combining all terms, we obtain:
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
Computing optimal variance bounds
Plugging $q^{*}_{\bar{u}\mid u}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{\bar{u}\mid u}))$. We obtain:
@@
Derivation of the zero-variance formula
We start from the general expression:
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
Now Plugging both optimal densities simultaneously:
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}^{2}(\mathbf{x}_{u})}$

### v2

@@ 4 Numerical Experiments
4.1 Analytical illustration of variance reduction
In this section, we illustrate the effectiveness of an importance sampling transformation in reducing the variance of $\eta_{u}$ estimators. As a benchmark example, we consider the classical Sobol’ g-function [41]:
@@ 4.1 Analytical illustration of variance reduction
where $X_{i}\sim\mathcal{U}[0,1]$ are independent inputs. This function is a standard benchmark in sensitivity analysis due to its nonlinear, non-monotonic structure and the availability of analytical Sobol’ indices. We consider the 3-dimensional case $(k=3)$ with $a_{i}=i$ and focus on the second-order index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$. For this part, one could in principle rely on kernel estimators [18, 11, 12], but since all quantities are analytically tractable, we rely on their closed forms and focus on the impact of optimal importance densities on variance reduction. The third input $X_{3}$ is treated as an uncontrolled variable $W$. We consider Case A in Section 3.3, where $u=\emptyset$. In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically. Specifically, the optimal density $q^{*}$ to estimate $\eta_{\{1,2\}}$ is given by
@@
Remark. Although our example involves an uncontrolled variable $W=X_{3}$, its particular structure satisfies $\phi^{2}(x_{u})\propto m^{2}(x_{u})$, so that the function $S_{A}(x_{u})$ simplifies to $S_{A}(x_{u})=c\,m^{4}(x_{u})$ for some constant $c>0$. As a result, $q_{u}^{*}(x_{u})$ retains the same analytical form as in the noise-free case, namely $q_{u}^{*}(x_{u})\propto m^{2}(x_{u})$. This coincidence is specific to the g-function and does not hold in general models.
We first plot the shape of $q^{*}$ over $0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
@@
Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
(a) Optimal density $q^{*}(x_{1},x_{2})$
@@
Figure 1: Left: optimal sampling density $q^{*}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{*}$.
Parametric optimisation
@@
where the shape parameters $\alpha_{i},\beta_{i}$ are positive real numbers, i.e., $\alpha_{i},\beta_{i}\in\mathbb{R}_{+}^{*}$ for $i=1,2$. We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform one at $(\alpha,\beta)=(1,1)$. The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimises the asymptotic variance, namely
@@
The optimal configuration is found numerically around $(\alpha,\beta)\approx(0.7,\,0.7)$. This yields a minimal variance of approximately $0.3368$, corresponding to a reduction of about $60\%$ compared to the uniform case $(\alpha=\beta=1)$. The optimal parameters are marked in red on the heatmap.
4.2 Sampling and estimation example
To move closer to a realistic setting, we now consider an estimation scenario where the target quantity $\eta_{\{1,2\}}$ is unknown and must be estimated from the observed data. We work with the same Sobol’ g-function and replace analytical quantities by Monte Carlo estimates while using the oracle $m(\mathbf{x})$ to isolate the impact of $q^{*}$.
We compare two strategies:
  • Reference case: samples drawn from the uniform input $p$. In this case, we estimate
@@ 4.2 Sampling and estimation example
  • Importance sampling case: samples are drawn from an empirical estimate of the optimal density $q^{*}\propto m^{2}(x_{1},x_{2})\,\mathbf{1}_{[0,1]^{2}}(x)$, as shown in the first part, denoted by $\widehat{q}_{\theta}$. The corresponding estimator reads
[…]
The Beta parametric family identified numerically in the first example can still be used as a simple baseline. In the following, however, we focus on a more flexible parametric approximation of $q^{*}$. Since $q^{*}(x_{1},x_{2})\propto m^{2}(x_{1},x_{2})$, its evaluation requires calls to the model $g(x_{1},x_{2},X_{3})$ to compute the conditional mean $m(x_{1},x_{2})$; consequently, the generation of samples approximately following $q^{*}$ depends on model evaluations.
A preliminary sample $(\mathbf{x}^{(i)})_{i=1}^{N_{\mathrm{fit}}}$ is obtained by a simple acceptance–rejection algorithm [19]:
$\mathbf{x}^{(i)}\sim\mathrm{Unif}([0,1]^{2}),\qquad u^{(i)}\sim\mathrm{Unif}(0,M),\qquad\text{accept if }u^{(i)}\leq m^{2}(\mathbf{x}^{(i)}),$
with $M=\sup_{[0,1]^{2}}m^{2}(\mathbf{x})$. The accepted points follow $q^{*}$ up to normalisation and are used to fit $q_{\theta}$.
Figure 3: Samples generated via rejection sampling following $q^{*}$.
To preserve the support $0,1]^{2}$ while retaining flexibility, we use a logit–Gaussian mixture model (logit–GMM), where $\mathrm{logit}(x)=\log\!\tfrac{x}{1-x}$ and $\mathrm{logistic}(z)=(1+e^{-z})^{-1}$. Each pilot point $\mathbf{x}=(x_{1},x_{2})$ is mapped to the unbounded space $\mathbf{z}=(z_{1},z_{2})$ through $z_{j}=\mathrm{logit}(x_{j})$, and a Gaussian mixture is then fitted:
$q_{Z}(\mathbf{z};\theta)=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(\mathbf{z};\mu_{k},\Sigma_{k}),\qquad\pi_{k}>0,\quad\sum_{k=1}^{K}\pi_{k}=1,$
with parameters $\theta=\{\pi_{k},\mu_{k},\Sigma_{k}\}_{k=1}^{K}$.
By the change-of-variables formula, this induces a properly normalised expression on $0,1]^{2}$:
$q_{\theta}(\mathbf{x})=q_{Z}\!\bigl(\mathrm{logit}(\mathbf{x});\theta\bigr)\prod_{j=1}^{2}\frac{1}{x_{j}(1-x_{j})},\qquad\mathbf{x}\in(0,1)^{2}.$
[…]
This experiment mimics a practical learning-based IS setting, where the optimal distribution $q^{*}$ is unknown and approximated from data, leading to noticeable variance reduction.
@@ 4.3 Distributional Sensitivity via Reverse Importance Sampling
We study the sensitivity of the quantity $\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m^{2}(\mathbf{X}_{u})\right],$ with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$. This falls into the framework of reverse importance sampling. The rank estimator [22], based on the sequential correlation of outputs, takes as input a sequence $Y^{(i)}$ sorted with respect to the associated inputs $\mathbf{X}_{u}^{(i)}$, and converges to $\eta_{u}$ via the average of consecutive products. By using the results from Section 3.1, the same estimator can be extended to compute $\eta_{u}(\theta)$ under any distribution $p_{\theta}$. This formulation allows extending the rank-based estimator of [22] to any perturbed distribution $p_{\theta}$ through reverse importance sampling and remains consistent. Its asymptotic behavior follows directly from the consistency result established in Proposition 3.2 of [22]. We note that in this section we work with the rank estimator (and not the lagged–rank estimator), which is not asymptotically efficient in terms of variance. Since our purpose here is only to study the sensitivity of the Sobol’ indices under different input distributions, the optimal densities derived in Section 3 are not involved. For completeness, we nevertheless derive in Appendix B the optimal sampling densities associated with the rank estimator itself. For each $\theta$, we define the reweighted output and the rank-based estimator becomes
$\widehat{\eta}^{\mathrm{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)},\qquad Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\,\frac{p_{\theta,-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})}{p_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})}\,Y.$ (20)
@@
We consider a black-box simulator based on the IGLOO2D code developed at ONERA [48], which returns the ice-accretion thickness along an airfoil profile. Each run of the simulator takes as input a vector of four physical variables $\mathbf{X}=(X_{1},X_{2},X_{3},X_{4})\in\mathbb{R}^{4}$, initially sampled independently from uniform distributions defined over the intervals $\mathbf{x}_{\mathrm{low}}=(3.5,\;65,\;261.15,\;4\times 10^{-4})$ and $\mathbf{x}_{\mathrm{up}}=(4.5,\;75,\;265.15,\;6\times 10^{-4})$, which respectively correspond to $X_{1}$ (angle of attack, AoA), $X_{2}$ (freestream velocity, $U_{\infty}$), $X_{3}$ (ambient temperature, $T_{\mathrm{inf}}$), and $X_{4}$ (liquid water content, LWC). A design of experiments of size $N=2500$ has been generated, yielding input samples $\mathbf{X}^{(i)}\in\mathbb{R}^{4}$ for $i=1,\dots,N$, and corresponding outputs $Y^{(i)}\in\mathbb{R}$, which represent the accretion thickness at a given location along the airfoil. All input variables $X_{j}$ are first normalised to the unit interval via $X_{j}^{\mathrm{std}}:=(X_{j}-x_{\mathrm{low},j})/(x_{\mathrm{up},j}-x_{\mathrm{low},j})$. To explore the sensitivity of this output to changes in the input distributions, we consider alternative input distributions $p_{\theta_{j}}(X_{j})$ built from rescaled $\operatorname{Beta}(\theta_{j})$ marginals on the same support, where varying $\theta_{j}=(\alpha_{j},\beta_{j})$ allows smooth deviations from the uniform case $(1,1)$. For any configuration $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$, the corresponding Sobol’ index $S_{u}(\boldsymbol{\theta})$ or contrast $\eta_{u}(\boldsymbol{\theta})$ can be estimated from the same dataset by importance weighting, using the factor
@@
With the rank estimator in Eq. (20), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$. Two perturbation strategies are considered: (i) marginal perturbation, where only the pair $\theta_{j}=(\alpha_{j},\beta_{j})$ associated with a single input $X_{j}$ is varied while the remaining marginals remain uniform; and (ii) global perturbation, where several or all input marginals are varied simultaneously. To keep the presentation simple, we focus on marginal perturbations. Specifically, we assess the distributional sensitivity of the contrast $\eta_{u}$ with $u=\{4\}$ by varying $\theta_{j}$ for one input variable at a time, keeping all others unchanged. This yields a family of perturbed estimates $\eta_{4}(\theta_{j})$, each revealing how the index $\eta_{4}$ responds to changes in the distribution of a specific input $X_{j}$, for $j=1,\dots,4$. All computations are based on the same reference dataset.
@@
These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, which varies through the two parameters $\alpha_{j}$ and $\beta_{j}$ of the Beta distribution, with different behaviours across the four inputs. The strongest variations are observed for the pair $(\alpha_{4},\beta_{4})$, which is expected since $\eta_{4}$ quantifies the influence of $X_{4}$. Overall, these results confirm that the proposed given‐data IS framework offers a powerful and cost-effective means to perform detailed distributional local sensitivity analyses of Sobol’ indices.
@@ 5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution that is different from the reference one. In particular, we show that by choosing suitable auxiliary densities, the asymptotic variance of efficient estimators can be further reduced, as changing the sampling distribution modifies the efficiency bound itself. In some cases, zero variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. In realistic settings where model evaluations are available, one may nevertheless approximate these optimal densities, which naturally leads to significant variance reduction. We also show that combining our method with the rank-based estimator [22] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [32]. The first limitation of our present work is that the optimal densities depend on the code, so additional computational effort is required. Another limitation is that we have proved the existence of one optimal distribution for every Sobol’ index, so it may be computationally expensive to estimate all these optimal densities, which increases the numerical effort. In a related direction, [15] investigate how several expectations under different target distributions can be estimated from a single Monte Carlo sample, using an adaptive importance sampling scheme with shared control variates. This suggests that multi–target designs could also be explored to construct sampling densities efficient for several Sobol’ quantities simultaneously. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
Acknowledgement
We gratefully acknowledge our ONERA colleagues — Lokman Bennani, Ghislain Blanchard, Maxime Bouyges, and Patricia Klotz — for providing the ice-accretion dataset. We also thank the reviewers for their valuable remarks during the revision process.
Supplementary Materials
All simulation codes for the academic experiments (Section 4) are available in the Supplementary Materials and on GitHub:
$\text{{https://github.com/HaythemBC/Importance-sampling-for-Sobol-indices-estimation}}.$
Data Availability Statement
All data used in the academic case studies are generated by simulation and can be fully reproduced using the associated code. The real-world dataset of Section 4.3 is confidential and cannot be shared publicly, but can be provided to reviewers upon reasonable request, in accordance with the data-sharing agreement of the contributing institution.
Appendix A Lemma 3.1 and other derivations
Proof of Lemma 3.1. We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
@@ Acknowledgement
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m(\mathbf{X}_{u}),$
@@
$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m^{2}(\mathbf{X}_{u})]$
This completes the proof.
Derivation of the optimal variance under $p$ (Eq. (12)). From Lemma 3.1, we recall
$\mathbb{E}_{q}[Z_{u}\mid X_{u}]=\sqrt{w_{u}(X_{u})}\,m(X_{u}),\qquad\bigl(\mathbb{E}_{q}[Z_{u}\mid X_{u}]\bigr)^{2}=w_{u}(X_{u})m^{2}(X_{u}).$
$\nu_{q}(X_{u}):=\mathbb{E}_{p}[w_{-u\mid u}\,\phi^{2}(X)\mid X_{u}]-m^{2}(X_{u}).$
We now compute the conditional second moment. Using the definition of $Z_{u}$ and $\phi^{2}(x)$, we obtain
$\mathbb{E}_{q}[Z_{u}^{2}\mid X_{u}]$ $=\mathbb{E}_{q}\!\left[w_{u}(X_{u})w_{-u\mid u}(X_{-u}\mid X_{u})Y^{2}\mid X_{u}\right]$
$=w_{u}(X_{u})\,\mathbb{E}_{p}\!\left[w_{-u\mid u}(X_{-u}\mid X_{u})\,\phi^{2}(X)\mid X_{u}\right].$
Hence
$\operatorname{Var}_{q}(Z_{u}\mid X_{u})=w_{u}(X_{u})\Big(\mathbb{E}_{p}[w_{-u\mid u}\,\phi^{2}(X)\mid X_{u}]-m^{2}(X_{u})\Big)=w_{u}(X_{u})\,\nu_{q}(X_{u}),$
Substituting all expressions into Proposition 3.2 yields
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=\mathbb{E}_{q}\!\left[4\,w_{u}(X_{u})m^{2}(X_{u})\,w_{u}(X_{u})\nu_{q}(X_{u})+w_{u}^{2}(X_{u})m^{4}(X_{u})\right]-\eta_{u}^{2}$
$=\mathbb{E}_{q}\!\left[w_{u}^{2}(X_{u})\bigl(4m^{2}(X_{u})\nu_{q}(X_{u})+m^{4}(X_{u})\bigr)\right]-\eta_{u}^{2}.$
Using the usual identity $\mathbb{E}_{q}[w_{u}^{2}g]=\mathbb{E}_{p}[w_{u}g]$, we finally obtain
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\!\left[w_{u}(X_{u})\bigl(4m^{2}(X_{u})\nu_{q}(X_{u})+m^{4}(X_{u})\bigr)\right]-\eta_{u}^{2},$
which is Eq. (12) of the main text.
Computing optimal variance bounds. Plugging $q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{-u\mid u}))$. We obtain:
@@
Derivation of the zero variance formula We start from the general expression:
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{-u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
Now plugging both optimal densities simultaneously.
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{-u}^{*}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u}(\mathbf{x}_{-u})\,f(\mathbf{x})}{m^{2}(\mathbf{x}_{u})}$